# Resource-Efficient Beam Prediction in mmWave Communications with Multimodal Realistic Simulation Framework

### v1

@@ Abstract
Beamforming is a key technology in millimeter-wave (mmWave) communications that improves signal transmission by optimizing directionality and intensity. However, conventional channel estimation methods, such as pilot signals or beam sweeping, often fail to adapt to rapidly changing communication environments. To address this limitation, multimodal sensing-aided beam prediction has gained significant attention, using various sensing data from devices such as LiDAR, radar, GPS, and RGB images to predict user locations or network conditions. Despite its promising potential, the adoption of multimodal sensing-aided beam prediction is hindered by high computational complexity, high costs, and limited datasets. Thus, in this paper, a resource-efficient learning approach is proposed to transfer knowledge from a multimodal network to a monomodal (radar-only) network based on cross-modal relational knowledge distillation (CRKD), while reducing computational overhead and preserving predictive accuracy. To enable multimodal learning with realistic data, a novel multimodal simulation framework is developed while integrating sensor data generated from the autonomous driving simulator CARLA with MATLAB-based mmWave channel modeling, and reflecting real-world conditions. The proposed CRKD achieves its objective by distilling relational information across different feature spaces, which enhances beam prediction performance without relying on expensive sensor data. Simulation results demonstrate that CRKD efficiently distills multimodal knowledge, allowing a radar-only model to achieve $94.62\%$ of the teacher performance. In particular, this is achieved with just $10\%$ of the teacher network’s parameters, thereby significantly reducing computational complexity and dependence on multimodal sensor data.
@@ I Introduction and Background
Millimeter-wave (mmWave) communications are widely recognized as a key enabler for next-generation wireless systems, because of their ability to provide high data rates and support numerous bandwidth-intensive applications [2]. A cornerstone technology in mmWave communication is beamforming, which directs wireless signals to specific spatial directions to improve signal strength and transmission quality [3]. However, mmWave signals experience high path loss and narrow beamwidth. This, in turn, makes it challenging to perform accurate beam alignment to maintain reliable links, particularly in dynamic or high mobility scenarios. A commonly employed solution for beam alignment is beam sweeping, in which the transmitter (Tx) and receiver (Rx) systematically scan multiple beam directions to find the optimal alignment. Although this approach is straightforward to implement, it can be inefficient in rapidly changing environments, introducing considerable overhead in terms of time and energy [4].
One promising approach for overcoming these limitations is to leverage multimodal sensing-aided beam prediction using data from sensors such as LiDAR, radar, GPS, and RGB cameras to monitor user trajectories and network conditions [5]. By incorporating this information, multimodal systems can forego or accelerate sequential beam scanning, allowing quicker and more flexible beam alignment with reduced overhead [6]. This technique is well-suited for applications that require low latency, such as autonomous driving and drone communications [7]. As such, multimodal sensing-aided beam prediction is becoming an important technology for future wireless networks [8]. However, designing practical multimodal sensing-aided beam prediction approaches requires overcoming a number of key challenges. First, deploying high-resolution sensors such as LiDAR or high-frame rate cameras at every base station is costly, and extensive use of cameras also raises privacy concerns [9]. Second, large-scale transformer-based fusion models are often needed to effectively leverage sensor data, but they can be computationally heavy [10, 11]. Third, creating comprehensive multimodal datasets is non-trivial, as existing public datasets often lack certain sensing modalities or have environment-specific constraints.
@@
DeepSense 6G [12] Measurement
WLADO [13] Measurement
Vi-Fi [14] Measurement
NEU [6] Measurement
DeepMIMO [15] Simulation
LASSE [16] Simulation
ViWi [17] Simulation
V2X-Sim [18] Simulation
e-Flash [19] Simulation
@@ I-A Prior Works
There has been a number of works that attempted to address the aforementioned challenges [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 12, 13, 14, 15, 16, 17, 18, 19, 30, 31, 32], as detailed next.
@@ I-A1 Traditional Approaches for mmWave Beamforming:
Traditional beamforming in mmWave systems has largely relied on exhaustive beam sweeping or heuristic codebook-based methods [20]. Beam sweeping is the most established beam optimization technique in mmWave communications, in which the transmitter and receiver systematically scan multiple beam directions to determine the one that provides the highest signal strength [21]. Although straightforward to implement and effective in static or low-mobility environments, beam sweeping can lead to considerable overhead and latency, which becomes problematic in high-speed or rapidly varying channels where real-time adaptability is essential [7]. To mitigate this issue, heuristic methods have also been explored, such as codebook-based beam selection, where candidate beams are rapidly identified, and greedy approaches that pick the beam delivering the highest instantaneous signal strength [22]. Although relatively simple and intuitive, these methods do not guarantee a global optimum and often fail in highly dynamic environments [4]. Consequently, relying solely on beam sweeping or heuristic approaches to overcome frequent blockages and severe path loss in mmWave systems can lead to increased communication delays and increased energy consumption, ultimately hampering the reliability of networks that demand rapid and continuous beam adaptation [23].
@@ I-A2 Multimodal Sensing-aided Beam Prediction:
To address the challenges of mmWave beamforming, recent studies propose to take advantage of sensor modalities such as LiDAR, radar, GPS, and RGB cameras for more accurate beamforming [24, 25, 26, 27, 28, 29]. The works in [24] and [25] used multimodal learning approaches that integrate LiDAR, radar, RGB, and GPS data to improve beam prediction accuracy. In [26], the authors proposed a deep quantum transformer network that fuses multimodal sensing data for robust mmWave beam prediction in integrated sensing and communication systems, demonstrating notable performance gains in real-world V2I scenarios. However, practical deployment of the solutions in [24, 25, 26] is hindered by sensor cost, increased computational complexity, and scalability concerns. Many existing models [27, 28, 29] for beam prediction assume the availability of various multimodal data, which may not be feasible in real-world deployments where the infrastructure is restricted to limited sensor modalities. In particular, current LiDAR systems are often prohibitively expensive, and the widespread deployment of cameras raises significant privacy concerns. Moreover, while transformer architectures like the ones used in [24, 25, 26, 27] can handle numerous multimodal tasks, they typically require significant computational resources to achieve fast inference. Without such hardware, real-world base stations cannot integrate such transformer-based solutions for beamforming purposes.
@@ I-A3 Multimodal Datasets and Simulation Frameworks:
To effectively train multimodal beam prediction models, comprehensive datasets containing multiple sensor modalities along with corresponding beamforming information are required. Table I shows publicly available datasets for multimodal learning in wireless communications. Moreover, these datasets can be broadly categorized into real-world and virtual environment datasets, each with its own significant limitations. In particular, publicly available datasets often lack the comprehensive multimodal data necessary for effective learning and evaluation. Among real-world datasets [12, 13, 14, 6], very few comprehensively include LiDAR, radar, and RGB data together. Hence, such approaches have very limited applicability in multimodal learning and sensor fusion research. Furthermore, most datasets are designed with a specific research focus, limiting their adaptability to explore new communication paradigms beyond their original scope. In addition, virtual environment datasets [15, 16, 17, 18, 19], often generated using network simulation tools such as MATLAB or Sionna, differ significantly from real-world datasets in terms of environmental complexity, sensor characteristics, and noise modeling. These discrepancies reduce the generalizability of models trained solely on synthetic data, making them less effective in practical deployment scenarios.
@@
To overcome these challenges, recent studies [30, 31, 32] proposed a realistic multimodal simulation framework that closely replicates real-world conditions in a virtual environment by integrating sensor data with beamforming information. This framework bridges the gap between synthetic and real-world datasets, enabling more robust and generalizable multimodal learning models for beam prediction in next-generation communication systems. To enhance the accuracy and realism of the generated data, the realistic multimodal simulation framework integrates an autonomous driving simulation platform with network simulation tools, as illustrated in Fig. 1. The authors in [30] proposed a multimodal simulation framework for digital twin (DT) enabled vehicle-to-everything (V2X) communications, using CARLA for realistic sensor data generation and Remcom Wireless InSite for precise ray-tracing-based wireless channel modeling, demonstrated through a blockage handover task for V2X link restoration. Similarly, the work in [31] introduced M3SC, a comprehensive multimodal sensing-communication dataset, generated using AirSim, WaveFarer, and Wireless InSite, effectively aligning the physical space (LiDAR, RGB, radar) with the electromagnetic space (mmWave, channel impulse response (CIR) matrices) under various weather conditions and frequency bands. Furthermore, in [32], the authors developed MVX-ViT, a co-simulation framework that integrates CARLA and Sionna to generate a multimodal V2X dataset, enabling AI-driven antenna position optimization. However, prior work has largely focused on perception tasks and lacks comprehensive experimentation on beamforming communication. To address this limitation, we propose a realistic multimodal simulation framework that combines CARLA with MATLAB, enabling detailed and diverse experiments on multimodal sensing and mmWave beamforming communication.
@@ I-A4 Knowledge Distillation for Resource-Efficient Learning:
Knowledge distillation (KD) compresses a high-capacity teacher model into a lightweight student model by training the student to mimic the teacher’s output, thus preserving performance while significantly reducing computational requirements [33]. This approach is particularly beneficial in resource-constrained environments, such as edge devices or mobile platforms, where large models are impractical due to limited memory and power budgets. In standard KD, the teacher’s probabilistic outputs (soft labels) guide the student, allowing it to learn richer data distributions than is typically possible with hard labels alone. Beyond logit-based distillation, other KD variants include feature-based distillation, which transfers intermediate representations, and relation-based distillation, which preserves the structural relationships between data points in the latent space of the teacher [34]. Although most KD methods assume that both the teacher and the student share the same input modality, cross-modal knowledge distillation (CKD) extends KD to allow knowledge transfer between models trained in different sensor modalities [35]. In [36], the authors proposed a teacher model trained on LiDAR and RGB data that can transfer its learned representations to a student model using radar and RGB input, thereby reducing the reliance on computationally expensive sensors during inference. Similarly, other CKD approaches [37] and [38] have been applied primarily to perception tasks such as object detection or classification. Although these works [36, 37, 38] demonstrate the potential of CKD for efficient inference, it does not explore or address beamforming communication. Thus, its applicability to real-time, resource-constrained beam prediction in 6G networks remains limited. To address these challenges, we investigate a radar-only student model for beam prediction that benefits from cross-modal relational knowledge distilled from transformer based multimodal teacher.
@@ I-B Contributions
  • Cross-modal relational knowledge distillation: We introduce a new method that transfers relational features from a multimodal teacher model (trained with LiDAR, radar, GPS, and RGB) to a student with only radar. This approach preserves predictive accuracy while reducing sensor dependencies and computational complexity.
@@ II System model
As illustrated in Fig. 2, we consider multimodal sensing-based beam prediction for downlink mmWave communications in multiple V2I environment, which consist of a mmWave beamformer with a uniform rectangular array (URA) system of $N$ antennas and a set $\mathcal{V}$ of $V$ vehicles, each with a single antenna. The beamformer employs a predefined beam codebook =$\mathcal{B}=\{c_{1},\cdots,c_{B}\}$ of size $B$, where each $c_{b}$ corresponds to a beam pattern realized by a weight vector $\boldsymbol{w}_{b}\in\mathbb{C}^{N}$. The beamformer selects the next optimal future beam index using a deep neural network (DNN) based on the sensing data from the previous observation window of size $P$. The sensing information at the channel sampling interval $t\in\mathcal{T}$ can be written as $\mathcal{X}=\{x[t-P+1],\cdots,x[t]\}$. We also consider two types of beamformers with different sensing configurations:
@@ II-A Network model
When a beamforming weight vector $\boldsymbol{w}_{b}$ of a beam pattern $c_{b}\in\mathcal{B}$ is applied, the magnitude of the response of the array for a vehicle $v$ over the path $l$ will be given by:
$R^{\mathrm{tx}}_{v,l}(c_{b})=10\log_{10}|\boldsymbol{w}_{b}\cdot\boldsymbol{a}_{v}(\theta_{l},\phi_{l})|,$ (1)
where $\boldsymbol{a}_{v}(\theta_{l},\phi_{l})$ is the array response at azimuth angle $\theta_{l}$, and elevation angle $\phi_{l}$. Accordingly, the total received signal strength (RSS) for vehicle $v$ using beam pattern $c_{b}$ is determined by summing the transmit power contributions of all $L$ paths and subtracting the respective path losses as follows:
$S_{v}(c_{b})=\sum_{l=1}^{L}\Bigl{(}R^{\mathrm{tx}}_{v,l}(c_{b})-P^{l}_{v}\Bigr{)},$ (2)
where $R^{\mathrm{tx}}_{v,l}(c_{b})$ is given by (1), and $P^{l}_{v}$ is the path loss of vehicle $v$ over path $l$. This path loss term typically accounts for distance-dependent attenuation, atmospheric absorption, and shadowing. A commonly used path loss model can be expressed as:
$P_{v}^{l}(\mathrm{dB})=P_{0}+10\alpha\log_{10}\!\bigl{(}d_{v,l}\bigr{)}+\chi_{\sigma},$ (3)
where $P_{0}$ is a reference path loss at a unit distance (e.g., 1 m), $\alpha$ is the path loss exponent, $d_{v,l}$ is the distance of the $l$-th path, and $\chi_{\sigma}$ captures large-scale fading effects such as shadowing or blockage.

### v2

@@ Abstract
Beamforming is a key technology in millimeter-wave (mmWave) communications that improves signal transmission by optimizing directionality and intensity. However, conventional channel estimation methods, such as pilot signals or beam sweeping, often fail to adapt to rapidly changing communication environments. To address this limitation, multimodal sensing-aided beam prediction has gained significant attention, using various sensing data from devices such as LiDAR, radar, GPS, and RGB images to predict user locations or network conditions. Despite its promising potential, the adoption of multimodal sensing-aided beam prediction is hindered by high computational complexity, high costs, and limited datasets. Thus, in this paper, a novel resource-efficient learning framework is introduced for beam prediction, which leverages a custom-designed cross-modal relational knowledge distillation (CRKD) algorithm specifically tailored for beam prediction tasks, to transfer knowledge from a multimodal network to a radar-only student model, achieving high accuracy with reduced computational cost. To enable multimodal learning with realistic data, a novel multimodal simulation framework is developed while integrating sensor data generated from the autonomous driving simulator CARLA with MATLAB-based mmWave channel modeling, and reflecting real-world conditions. The proposed CRKD achieves its objective by distilling relational information across different feature spaces, which enhances beam prediction performance without relying on expensive sensor data. Simulation results demonstrate that CRKD efficiently distills multimodal knowledge, allowing a radar-only model to achieve $94.62\%$ of the teacher performance. In particular, this is achieved with just $10\%$ of the teacher network’s parameters, thereby significantly reducing computational complexity and dependence on multimodal sensor data.
@@ I Introduction and Background
Millimeter-wave (mmWave) communications are widely recognized as a key enabler for next-generation wireless systems, because of their ability to provide high data rates and support numerous bandwidth-intensive applications [tataria20216g]. A cornerstone technology in mmWave communication is beamforming, which directs wireless signals to specific spatial directions to improve signal strength and transmission quality [brilhante2023literature]. However, mmWave signals experience high path loss and narrow beamwidth. This, in turn, makes it challenging to perform accurate beam alignment to maintain reliable links, particularly in dynamic or high mobility scenarios. A commonly employed solution for beam alignment is beam sweeping, in which the transmitter (Tx) and receiver (Rx) systematically scan multiple beam directions to find the optimal alignment. Although this approach is straightforward to implement, it can be inefficient in rapidly changing environments, introducing considerable overhead in terms of time and energy [li2020beam].
One promising approach for overcoming these limitations is to leverage multimodal sensing-aided beam prediction using data from sensors such as LiDAR, radar, GPS, and RGB cameras to monitor user trajectories and network conditions [wen2024survey]. By incorporating this information, multimodal systems can forego or accelerate sequential beam scanning, allowing quicker and more flexible beam alignment with reduced overhead [salehi2022deep]. This technique is well-suited for applications that require low latency, such as autonomous driving and drone communications [xiao2021survey]. As such, multimodal sensing-aided beam prediction is becoming an important technology for future wireless networks [khan2023machine]. However, designing practical multimodal sensing-aided beam prediction approaches requires overcoming a number of key challenges. First, deploying high-resolution sensors such as LiDAR or high-frame rate cameras at every base station is costly, and extensive use of cameras also raises privacy concerns [jiang2022lidar]. Second, large-scale transformer-based fusion models are often needed to effectively leverage sensor data, but they can be computationally heavy [wu2022blockage, cheng2023intelligent]. Third, creating comprehensive multimodal datasets is non-trivial, as existing public datasets often lack certain sensing modalities or have environment-specific constraints.
@@
DeepSense 6G [alkhateeb2023deepsense] Measurement
WLADO [waldo_dataset] Measurement
Vi-Fi [liu2022vi] Measurement
NEU [salehi2022deep] Measurement
DeepMIMO [Alkhateeb2019] Simulation
LASSE [klautau20185g] Simulation
ViWi [Alrabeiah19] Simulation
V2X-Sim [li2022v2x] Simulation
e-Flash [steinmetzer2017compressive] Simulation
@@ I-A Prior Works
There has been a number of works that attempted to address the aforementioned challenges [kutty2015beamforming, heng2021six, que2023joint, zhang2017codebook, alrabeiah2022neural, bian20243, nie2023vision, cui2024sensing, tariq2024deep, tian2023multimodal, charan2022vision, zhu2025advancing, alkhateeb2023deepsense, waldo_dataset, liu2022vi, Alkhateeb2019, klautau20185g, Alrabeiah19, li2022v2x, steinmetzer2017compressive, cazzella2024multi, cheng2023m, gharsallah2024mvx], as detailed next.
@@ I-A1 Traditional Approaches for mmWave Beamforming:
Traditional beamforming in mmWave systems has largely relied on exhaustive beam sweeping or heuristic codebook-based methods [kutty2015beamforming]. Beam sweeping is the most established beam optimization technique in mmWave communications, in which the transmitter and receiver systematically scan multiple beam directions to determine the one that provides the highest signal strength [heng2021six]. In [que2023joint], the authors leveraged simultaneous localization and mapping (SLAM) techniques to significantly improve beam tracking and management in mmWave systems by incorporating geometric consistency and environmental features. Although straightforward to implement and effective in static or low-mobility environments, beam sweeping can lead to considerable overhead and latency, which becomes problematic in high-speed or rapidly varying channels where real-time adaptability is essential [xiao2021survey]. To mitigate this issue, heuristic methods have also been explored, such as codebook-based beam selection, where candidate beams are rapidly identified, and greedy approaches that pick the beam delivering the highest instantaneous signal strength [zhang2017codebook]. Although relatively simple and intuitive, these methods do not guarantee a global optimum and often fail in highly dynamic environments [xiao2021survey]. Consequently, relying solely on beam sweeping or heuristic approaches to overcome frequent blockages and severe path loss in mmWave systems can lead to increased communication delays and increased energy consumption, ultimately hampering the reliability of networks that demand rapid and continuous beam adaptation [alrabeiah2022neural].
@@ I-A2 Multimodal Sensing-aided Beam Prediction:
To address the challenges of mmWave beamforming, recent studies propose to take advantage of sensor modalities such as LiDAR, radar, GPS, and RGB cameras for more accurate beamforming [bian20243, nie2023vision, cui2024sensing, tariq2024deep, tian2023multimodal, charan2022vision, zhu2025advancing]. The works in [nie2023vision, bian20243, cui2024sensing] used multimodal learning approaches that integrate LiDAR, radar, RGB, and GPS data to improve beam prediction accuracy. In [tariq2024deep], the authors proposed a deep quantum transformer network that fuses multimodal sensing data for robust mmWave beam prediction in integrated sensing and communication systems, demonstrating notable performance gains in real-world V2I scenarios. However, practical deployment of the solutions in [nie2023vision, cui2024sensing, tariq2024deep] is hindered by sensor cost, increased computational complexity, and scalability concerns. Many existing models [tian2023multimodal, charan2022vision, zhu2025advancing] for beam prediction assume the availability of various multimodal data, which may not be feasible in real-world deployments where the infrastructure is restricted to limited sensor modalities. In particular, current LiDAR systems are often prohibitively expensive, and the widespread deployment of cameras raises significant privacy concerns. Moreover, while transformer architectures like the ones used in [nie2023vision, cui2024sensing, tariq2024deep, tian2023multimodal] can handle numerous multimodal tasks, they typically require significant computational resources to achieve fast inference. Without such hardware, real-world base stations cannot integrate such transformer-based solutions for beamforming purposes.
@@ I-A3 Multimodal Datasets and Simulation Frameworks:
To effectively train multimodal beam prediction models, comprehensive datasets containing multiple sensor modalities along with corresponding beamforming information are required. Table I shows publicly available datasets for multimodal learning in wireless communications. Moreover, these datasets can be broadly categorized into real-world and virtual environment datasets, each with its own significant limitations. In particular, publicly available datasets often lack the comprehensive multimodal data necessary for effective learning and evaluation. Among real-world datasets [alkhateeb2023deepsense, waldo_dataset, liu2022vi, salehi2022deep], very few comprehensively include LiDAR, radar, and RGB data together. Hence, such approaches have very limited applicability in multimodal learning and sensor fusion research. Furthermore, most datasets are designed with a specific research focus, limiting their adaptability to explore new communication paradigms beyond their original scope. In addition, virtual environment datasets [Alkhateeb2019, klautau20185g, Alrabeiah19, li2022v2x, steinmetzer2017compressive], often generated using network simulation tools such as MATLAB or Sionna, differ significantly from real-world datasets in terms of environmental complexity, sensor characteristics, and noise modeling. These discrepancies reduce the generalizability of models trained solely on synthetic data, making them less effective in practical deployment scenarios.
@@
To overcome these challenges, recent studies [cazzella2024multi, cheng2023m, gharsallah2024mvx] proposed a realistic multimodal simulation framework that closely replicates real-world conditions in a virtual environment by integrating sensor data with beamforming information. This framework bridges the gap between synthetic and real-world datasets, enabling more robust and generalizable multimodal learning models for beam prediction in next-generation communication systems. To enhance the accuracy and realism of the generated data, the realistic multimodal simulation framework integrates an autonomous driving simulation platform with network simulation tools, as illustrated in Fig. 1. The authors in [cazzella2024multi] proposed a multimodal simulation framework for digital twin (DT) enabled vehicle-to-everything (V2X) communications, using CARLA for realistic sensor data generation and Remcom Wireless InSite for precise ray-tracing-based wireless channel modeling, demonstrated through a blockage handover task for V2X link restoration. Similarly, the work in [cheng2023m] introduced M3SC, a comprehensive multimodal sensing-communication dataset, generated using AirSim, WaveFarer, and Wireless InSite, effectively aligning the physical space (LiDAR, RGB, radar) with the electromagnetic space (mmWave, channel impulse response (CIR) matrices) under various weather conditions and frequency bands. Furthermore, in [gharsallah2024mvx], the authors developed MVX-ViT, a co-simulation framework that integrates CARLA and Sionna to generate a multimodal V2X dataset, enabling AI-driven antenna position optimization. However, prior work has largely focused on perception tasks and lacks comprehensive experimentation on beamforming communication. To address this limitation, we propose a realistic multimodal simulation framework that combines CARLA with MATLAB, enabling detailed and diverse experiments on multimodal sensing and mmWave beamforming communication.
@@ I-A4 Knowledge Distillation for Resource-Efficient Learning:
Knowledge distillation (KD) compresses a high-capacity teacher model into a lightweight student model by training the student to mimic the teacher’s output, thus preserving performance while significantly reducing computational requirements [hinton2015distilling]. This approach is particularly beneficial in resource-constrained environments, such as edge devices or mobile platforms, where large models are impractical due to limited memory and power budgets. In standard KD, the teacher’s probabilistic outputs (soft labels) guide the student, allowing it to learn richer data distributions than is typically possible with hard labels alone. Beyond logit-based distillation, other KD variants include feature-based distillation, which transfers intermediate representations, and relation-based distillation, which preserves the structural relationships between data points in the latent space of the teacher [gou2021knowledge]. Although most KD methods assume that both the teacher and the student share the same input modality, cross-modal knowledge distillation (CKD) extends KD to allow knowledge transfer between models trained in different sensor modalities [yang2023categories]. In [zhao2024crkd], the authors proposed a teacher model trained on LiDAR and RGB data that can transfer its learned representations to a student model using radar and RGB input, thereby reducing the reliance on computationally expensive sensors during inference. Similarly, other CKD approaches [huo2024c2kd] and [ni2022cross] have been applied primarily to perception tasks such as object detection or classification. Although these works [zhao2024crkd, huo2024c2kd, ni2022cross] demonstrate the potential of CKD for efficient inference, it does not explore or address beamforming communication. Thus, its applicability to real-time, resource-constrained beam prediction in 6G networks remains limited. To address these challenges, we investigate a radar-only student model for beam prediction that benefits from cross-modal relational knowledge distilled from transformer based multimodal teacher.
@@ I-B Contributions
  • Cross-modal relational knowledge distillation: We introduce a new beam prediction–specialized method that transfers relational features from a multimodal teacher model (trained with LiDAR, radar, GPS, and RGB) to a student with only radar. This approach preserves predictive accuracy while reducing sensor dependencies and computational complexity.
@@ II System model
As illustrated in Fig. 2, we consider multimodal sensing-based beam prediction for downlink mmWave communications in multiple V2I environments, which consists of a mmWave beamformer with a uniform rectangular array (URA) system of $N$ antennas and a set $\mathcal{V}$ of $V$ vehicles, each with a single antenna. The beamformer employs a predefined beam codebook $\mathcal{B}=\{c_{1},\cdots,c_{B}\}$ of size $B$, where each $c_{b}$ corresponds to a beam pattern realized by a weight vector $\boldsymbol{w}_{b}\in\mathbb{C}^{N}$. The beamformer selects the next optimal future beam index using a deep neural network (DNN) based on the sensing data from the previous observation window of size $P$. The sensing information at the channel sampling interval $t\in\mathcal{T}$ can be written as $\mathcal{X}=\{x[t-P+1],\cdots,x[t]\}$. We also consider two types of beamformers with different sensing configurations:
@@ II-A Network model
To model the characteristics of mmWave communication, we adopt a geometric ray-tracing-based channel framework that captures sparse multipath, directional propagation, and frequency-selective path loss. All path-level parameters $\{\beta_{v,l},\theta_{l},\phi_{l},\tau_{l}\}_{l=1}^{L}$ are generated using MATLAB’s raytrace function, which implements a 3D shoot–bounce–ray (SBR) solver consistent with 3GPP TR 38.901. This setup includes specular reflections, single-edge diffractions (up to order 2), and realistic urban materials (e.g., concrete, glass, metal), modeled in an urban microcell environment imported from CARLA.
When applying a beamforming weight vector $\boldsymbol{w}_{b}$ of a beam pattern $c_{b}\in\mathcal{B}$, the magnitude of the response of the array for a vehicle $v$ over the path $l$ is calculated as:
$R^{\mathrm{tx}}_{v,l}(c_{b})=10\log_{10}\left|\boldsymbol{w}_{b}^{\mathrm{H}}\cdot\boldsymbol{a}_{v}(\theta_{l},\phi_{l})\right|,$ (1)
where $\boldsymbol{a}_{v}(\theta_{l},\phi_{l})\in\mathbb{C}^{N}$ is the array response vector at azimuth $\theta_{l}$ and elevation $\phi_{l}$, and H denotes the Hermitian (conjugate transpose). This response reflects the directional gain of the array for a specific path.
The total received signal strength (RSS) for vehicle $v$ using beam pattern $c_{b}$ is determined by aggregating contributions from all $L$ propagation paths, weighted by path losses:
$S_{v}(c_{b})=\sum_{l=1}^{L}\left(R^{\mathrm{tx}}_{v,l}(c_{b})-P^{l}_{v}\right),$ (2)
where $P^{l}_{v}$ represents the path loss of the $l$-th path. This loss includes distance-based attenuation, material-dependent reflection loss, and large-scale fading. A common empirical model expresses this as follows:
$P_{v}^{l}(\mathrm{dB})=P_{0}+10\alpha\log_{10}\!\left(d_{v,l}\right)+\chi_{\sigma},$ (3)
where $P_{0}$ is the reference path loss at 1 meter, $\alpha$ is the path loss exponent, $d_{v,l}$ is the distance of the $l$-th path, and $\chi_{\sigma}\sim\mathcal{N}(0,\sigma^{2})$ models shadow effects.
In addition, the effective channel gain for each path includes a carrier-dependent phase shift due to propagation delay $\tau_{v,l}$, expressed as:
$h_{v,l}=\sqrt{\beta_{v,l}}\,e^{-j2\pi f_{c}\tau_{v,l}},$ (4)
which is embedded in $\beta_{v,l}$ during the simulation. This realistic channel model reflects the sparse, high-directivity nature of mmWave environments and forms the foundation for evaluating beam prediction accuracy under our proposed knowledge distillation framework. Importantly, this formulation ensures that beam prediction models are trained and evaluated under physically realistic and path-dependent conditions, allowing them to learn angular selectivity and NLoS behavior intrinsic to real-world mmWave propagation.